# Variational Pseudo Marginal Methods for Jet Reconstruction in Particle Physics

### v1

@@ Abstract
Reconstructing jets in particle physics, which provide vital insights into the properties and histories of subatomic particles produced in high-energy collisions, is a main problem in data analyses in collider physics. This intricate task deals with estimating the latent structure of a jet (binary tree) and involves parameters such as particle energy, momentum, and types. While Bayesian methods offer a natural approach for handling uncertainty and leveraging prior knowledge, they face significant challenges due to the super-exponential growth of potential jet topologies as the number of observed particles increases. To address this, we introduce a Combinatorial Sequential Monte Carlo approach for inferring jet latent structures. As a second contribution, we leverage the resulting estimator to develop a variational inference algorithm for parameter learning. Building on this, we introduce a variational family using a pseudo-marginal framework for a fully Bayesian treatment of all variables, unifying the generative model with the inference process. We illustrate our method’s effectiveness through experiments using data generated with a collider physics generative model, highlighting superior speed and accuracy across a range of tasks.
@@ 1 Introduction
1Columbia University. 2Barnard College, Columbia University. 3 Telefonica Research 4 University of Amsterdam. Correspondence to amoretti@cs.columbia.edu. * Authors contributed equally.
@@
Reconstructing jets in particle physics deals with estimating a high-quality hierarchical clustering. A comprehensive approach to this process also involves inference on model parameters, which could provide insights into our understanding of quantum chromodynamics (QCD), i.e. the theory of the strong interaction between quarks mediated by gluons. Hierarchical clustering forms a natural data representation of data generated by a Markov tree, and has been applied in a wide variety of settings such as entity resolution for knowledge-bases [Green et al., 2012, Vashishth et al., 2018], personalization [Zhang et al., 2014], and jet physics [Cacciari et al., 2008, Catani et al., 1993, Dokshitzer et al., 1997, Ellis and Soper, 1993]. Typically, work has focused on approximate methods for relatively large datasets [Bateni et al., 2017, Monath et al., 2019, Naumov et al., 2020, Dubey et al., 2014, Hu et al., 2015, Monath et al., 2020, Dubey et al., 2020, Monath et al., 2021]. However, there are relevant use cases for hierarchical clustering that require exact or high-quality approximations on small to medium-sized datasets [Greenberg et al., 2020, 2021]. This paper deals with one of these use cases: reconstructing the latent hierarchy of jets in particle physics. Within this context, Bayesian methods provide a natural approach for handling uncertainty, but the super-exponential scaling of the number of hierarchies with the size of the datasets presents significant difficulties, i.e. the number of topologies grows as $(2N-3)!!$!, with $N$ being the number of leaves. This super-exponential growth in the space of configurations makes brute force and exact methods intractable.
ProtonProtonLeaves = Observed Particles
@@
Figure 1: Jets as binary trees. Schematic representation of the production of a jet at CERN’s LHC. Incoming protons collide, producing two new particles (light blue). Each new particle undergoes a sequence of binary splittings until stable particles (solid blue) are produced and measured by a detector.
Jet physics.
@@
During high-energy particle collisions, such as those observed at the Large Hadron Collider (LHC) at CERN, collimated sprays of particles called jets are produced. The jet constituents are the observed final-state particles that hit the detector and are originated by a showering process (described by QCD) where an initial (unstable) particle goes through successive binary splittings. Intermediate (latent) particles can be identified as internal nodes of a hierarchical clustering and the final-state (observed) particles correspond to the leaves in a binary tree. Fig. 1 provides a schematic representation of this process. This results in several possible latent topologies corresponding to a set of leaves. This representation, first suggested in Louppe et al. [2019], connects jets physics with natural language processing (NLP) and biology.
Collider data analysis.
@@
A main problem in data analyses of collider physics deals with estimating the latent showering process (hierarchical clustering) of a jet, which is needed for subsequent tasks that aim to identify the creation of different types of sub-atomic particles. The final goal is to perform precision measurements to test the predictions of the Standard Model of Particle Physics and explore potential models of new physics particles, thereby advancing our comprehension of the universe’s fundamental constituents. The improved performance of deep learning jet classifiers [Butter et al., 2019] (for the initial state particles) over traditional clustering-based physics observables gives evidence of the limitedness of current clustering algorithms since the right clustering should be optimal for classification tasks. Thus, high-quality approximations in this context would be highly beneficial for data analyses in experimental particle physics.
Jet simulators.
@@
Currently, there are high-fidelity simulations for jets, such as Pythia [Sjostrand et al., 2006], Herwig [Bellm et al., 2016], and Sherpa [Gleisberg et al., 2009]. These simulators are grounded in QCD but make several approximations that introduce parametric modeling choices that are not predicted from the underlying theory, so they are commonly tuned to match the data. Many tasks in jet physics can be framed in probabilistic terms [Cranmer et al., 2021]. In particular, we consider the challenges of calculating the maximum likelihood hierarchy given a set of leaves (jet constituents), the posterior distribution of hierarchies, as well as estimating the marginal likelihood. Also, these quantities are relevant to tune the parameters of the simulators. While these formulations are helpful conceptually, they are not practical in current high-fidelity simulations for jets, given that the likelihood is typically intractable (they are implicit models). Thus, we consider Ginkgo [Cranmer, Kyle et al., 2021, Cranmer et al., 2019]: a semi-realistic generative model for jets with a tractable joint likelihood and captures essential ingredients of parton shower generators in full physics simulations. In particular, Ginkgo was designed to enable implementations of probabilistic programming, differentiable programming, dynamic programming and variational inference.Within the analogy between jets and NLP, Ginkgo can be considered as ground-truth parse trees with a known language model.
Jet clustering.
@@
For each jet produced at the LHC, there is an inference task on the latent hierarchy that typically involves 10 to 100 particles (leaves). Though this is a relatively small number of elements, exhaustive solutions are intractable, and current exact methods, e.g., Greenberg et al. [2020, 2021], have limited scalability. The industry standard uses agglomerative clustering techniques, which are greedy and based on heuristics Cacciari et al. [2008], Catani et al. [1993], Dokshitzer et al. [1997], Ellis and Soper [1993], typically finding low-quality hierarchical clusterings. Regarding likelihood-based clustering (applied to Ginkgo datasets), previous work Greenberg et al. [2020] introduced a classical data structure and dynamic programming algorithm (the cluster trellis) that exactly finds the marginal likelihood over the space of configurations and the maximum likelihood hierarchy. Also, an A* search algorithm combined with a trellis data structure that finds the exact maximum likelihood hierarchy was introduced in  Greenberg et al. [2021]. Finally,  Cranmer et al. [2022] pairs Ginkgo with the cluster trellis [Greenberg et al., 2020], to use the marginal likelihood to directly characterize the discrimination power of the optimal classifier [J. Stuart and Arnold, 1994, Cranmer and Plehn, 2007] as well as to compute the exact maximum likelihood estimate for the simulator’s parameters. While these works provide exact algorithms that extend the reach of brute force methods, they have an exponential space and time complexity, becoming intractable for datasets with as few as 15 leaves. For this reason, Greenberg et al. [2020, 2021] also provide approximate solutions at the cost of finding lower-quality hierarchies.
Bayesian inference.
@@
A recent body of research has melded variational inference (VI) and sequential search. These connections are realized through the development of a variational family for hidden Markov models, employing Sequential Monte Carlo (Smc) as the marginal likelihood estimator [Maddison et al., 2017, Naesseth et al., 2018, Le et al., 2018, Moretti et al., 2019, 2020, 2021]. Within the field of Bayesian phylogenetics (the study of evolutionary histories), various methods have been proposed for inference on tree structures. Common approaches include local search algorithms like random-walk Mcmc [Ronquist et al., 2012] and sequential search algorithms like Combinatorial Sequential Monte Carlo (Csmc[Bouchard-Côté et al., 2012, Wang et al., 2015]. Mcmc methods also handle model learning.  Dinh et al. [2017] proposes ppHmc which extends Hamiltonian Monte Carlo to phylogenies. Evaluating the likelihood term in Mcmc acceptance ratios can be challenging. As a workaround, particle Mcmc (Pmcmc) algorithms use Smc to estimate the marginal likelihood and define Mcmc proposals for parameter learning [Wang and Wang, 2020].
@@
Pseudo-marginal methods are a class of statistical techniques used to approximate difficult-to-compute probabilities, typically by introducing auxiliary random variables to form an unbiased estimate of the target probability [Andrieu and Roberts, 2009].  Beaumont [2003] introduced a method in genetics to sample genealogies in a fully Bayesian framework. Tran et al. [2016] utilizes pseudo-marginal methods to perform variational Bayesian inference with an intractable likelihood. Our work is a synthesis of Wang et al. [2015] and Moretti et al. [2021] in that we introduce a variational approximation on topologies using Smc and a VI framework to learn parameters.
Contributions of this Paper:
@@
    We expand upon the Csmc technique introduced by Wang et al. [2015] and the Ncsmc method from Moretti et al. [2021] to introduce a Combinatorial Sequential Monte Carlo framework for inferring hierarchical clusterings for jets. The resulting estimators are unbiased and consistent. To the best of our knowledge, this marks the first adaptation of Smc methods to jet reconstruction in particle physics.
@@
    We leverage the resulting Smc estimators to develop two approximate posteriors on jet hierarchies and correspondingly two VI methods for parameter learning. We illustrate the effectiveness of both methods through experiments using data generated with Ginkgo [Cranmer, Kyle et al., 2021], highlighting superior speed and accuracy across various tasks.
@@
    We show how partial states and re-sampled indices generated by Smc can be interpreted as auxiliary random variables within a pseudo-marginal framework, thus establishing connections between variational pseudo-marginal methods and Vsmc [Naesseth et al., 2018, Moretti et al., 2021].
@@ 2 Background
Section 2.1 provides an overview of the Ginkgo generative model for jet physics [Cranmer, Kyle et al., 2021]. Section 2.2 summarizes the approximate inference techniques this work builds upon.
@@ 2.1.1 Generative process
In Ginkgo the generative process begins with the splitting of a parent (root) node with invariant mass squared $t_{P}$ into two children, as shown schematically in Fig. 2 (left). The process is characterized by a cutoff mass squared $t_{cut}$, and a rate parameter $\lambda$ for the exponential distribution governing the decay. During generation, as long as the invariant mass squared of a node exceeds the cutoff value ($t_{P}>t_{cut}$), that node is promoted to be a parent and the algorithm recursively splits it. The squared masses of each new left (L) and right (R) child nodes ($t_{L}$, $t_{R}$) are obtained from sampling from the exponential distribution $f(t|\lambda,t_{P})$ defined in Eq. 1, with parameters specific to each child ($L$, $R$). Finally, once $t_{L}$ and $t_{R}$ are sampled, the corresponding energy-momentum vectors ($z_{L}$, $z_{R}$) for the ($L$, $R$) nodes are derived from $t_{P}$, $t_{L}$ and $t_{R}$ following energy-momentum conservation rules, i.e. applying a 2-body particle decay (see  [Cranmer, Kyle et al., 2021, Cranmer et al., 2019] for more details). Next, we specify the exponential distribution
@@
$t_{P}=t(z_{P})>t_{cut}$$t_{L}\sim f(\cdot|\lambda,t_{P})$$t_{R}\sim f(\cdot|\lambda,(\sqrt{t_{P}}-\sqrt{t_{L}})^{2})$BC
@@
[…]
@@
Figure 2: Illustration of the Ginkgo generative and reconstruction processes. (a) Ginkgo starts with a parent node characterized by a 4-vector $z_{p}=(E,\vec{p})$ and invariant mass squared $t_{P}=E^{2}-|\vec{p}|^{2}$. If $t_{P}$ is greater than the cut off value $t_{cut}$, then the parent node splits (we have a particle decay). The left and right nodes invariant mass squared ($t_{L}$, $t_{R}$) are sampled from a truncated exponential distribution defined in Eq. 1. (b) The splitting likelihood reconstruction process of a node, defined in Eq. 4 begins with two child nodes $B$ and $C$ along with their respective 4-vectors $z_{L}$ and $z_{R}$. The 4-vector for the parent node $A$ is calculated as $z_{P}=z_{L}+z_{R}$ and then $t_{P}=t(z_{P})$. Next, we obtain $t_{L}=t(z_{L})$, $t_{R}=t(z_{R})$ and define $t_{P}^{L}=t_{P}$, and $t_{P}^{R}=(\sqrt{t_{P}}-\sqrt{t_{L}})^{2}$. Finally, the left splitting term likelihood $\ell(t_{L},\lambda,t_{cut},t_{P}^{L},t_{P})$ and the right one $\ell(t_{R},\lambda,t_{cut},t_{P}^{R},t_{P})$ are evaluated.

### v2

@@ Abstract
Reconstructing jets, which provide vital insights into the properties and histories of subatomic particles produced in high-energy collisions, is a main problem in data analyses of collider physics. This intricate task deals with estimating the latent structure of a jet (binary tree) and involves parameters such as particle energy, momentum, and types. While Bayesian methods offer a natural approach for handling uncertainty and leveraging prior knowledge, they face significant challenges due to the super-exponential growth of potential jet topologies as the number of observed particles increases. To address this, we introduce a Combinatorial Sequential Monte Carlo approach for inferring jet latent structures. As a second contribution, we leverage the resulting estimator to develop a variational inference algorithm for parameter learning. Building on this, we introduce a variational family using a pseudo-marginal framework for a fully Bayesian treatment of all variables, unifying the generative model with the inference process. We illustrate our method’s effectiveness through experiments using data generated with a collider physics generative model, highlighting superior speed and accuracy across a range of tasks.
@@ 1 Introduction
Reconstructing jets in particle physics deals with estimating a high-quality hierarchical clustering. A comprehensive approach to this process also involves inference on model parameters, which could provide insights into our understanding of quantum chromodynamics (QCD), i.e. the theory of the strong interaction between quarks mediated by gluons. Hierarchical clustering forms a natural data representation of data generated by a Markov tree, and has been applied in a wide variety of settings such as entity resolution for knowledge-bases (Green et al., 2012; Vashishth et al., 2018), personalization (Zhang et al., 2014), and jet physics (Cacciari et al., 2008; Catani et al., 1993; Dokshitzer et al., 1997; Ellis & Soper, 1993). Typically, work has focused on approximate methods for relatively large datasets (Bateni et al., 2017; Monath et al., 2019; Naumov et al., 2020; Dubey et al., 2014; Hu et al., 2015; Monath et al., 2020; Dubey et al., 2020; Monath et al., 2021). However, there are relevant use cases for hierarchical clustering that require exact or high-quality approximations on small to medium-sized datasets (Greenberg et al., 2020; 2021). This paper deals with one of these use cases: reconstructing the latent hierarchy of jets in particle physics. Within this context, Bayesian methods provide a natural approach for handling uncertainty, but the super-exponential scaling of the number of hierarchies with the size of the datasets presents significant difficulties, i.e. the number of topologies grows as $(2N-3)!!$!, with $N$ being the number of leaves. This super-exponential growth in the space of configurations makes brute force and exact methods intractable.
Jet generationJet reconstructionProtonProton
$\Longrightarrow$
Leaves = Observed Particles
@@
Figure 1: Jets as binary trees. Left: Schematic representation of the production of a jet at CERN’s LHC. Incoming protons collide, producing two new particles (light blue). Each new particle undergoes a sequence of binary splittings until stable particles (solid blue) are produced and measured by a detector. Right: In jet reconstruction, only the leaf nodes are observed and the tree topology is inferred. Each latent tree topology represents a different possible splitting history.
1.1 Jet physics
@@
During high-energy particle collisions, such as those observed at the Large Hadron Collider (LHC) at CERN, collimated sprays of particles called jets are produced. The jet constituents are the observed final-state particles that hit the detector and are originated by a showering process (described by QCD) where an initial (unstable) particle goes through successive binary splittings. Intermediate (latent) particles can be identified as internal nodes of a hierarchical clustering and the final-state (observed) particles correspond to the leaves in a binary tree. Fig. 1 provides a schematic representation of this process. This results in several possible latent topologies corresponding to a set of leaves. This representation, first suggested in Louppe et al. (2019), connects jets physics with natural language processing (NLP) and biology.
1.2 Collider data analysis
@@
A main problem in data analyses of collider physics deals with estimating the latent showering process (hierarchical clustering) of a jet, which is needed for subsequent tasks that aim to identify the creation of different types of sub-atomic particles. The final goal is to perform precision measurements to test the predictions of the Standard Model of Particle Physics and explore potential models of new physics particles, thereby advancing our comprehension of the universe’s fundamental constituents. The improved performance of deep learning jet classifiers (Butter et al., 2019) (for the initial state particles) over traditional clustering-based physics observables gives evidence of the limitedness of current clustering algorithms since the right clustering should be optimal for classification tasks. Thus, high-quality approximations in this context would be highly beneficial for data analyses in experimental particle physics.
1.3 Jet simulators
@@
Currently, there are high-fidelity simulations for jets, such as Pythia (Sjostrand et al., 2006), Herwig (Bellm et al., 2016), and Sherpa (Gleisberg et al., 2009). These simulators are grounded in QCD but make several approximations that introduce parametric modeling choices that are not predicted from the underlying theory, so they are commonly tuned to match the data. Many tasks in jet physics can be framed in probabilistic terms (Cranmer et al., 2021). In particular, we consider the challenges of calculating the maximum likelihood hierarchy given a set of leaves (jet constituents), the posterior distribution of hierarchies, as well as estimating the marginal likelihood. Also, these quantities are relevant to tune the parameters of the simulators. While these formulations are helpful conceptually, they are not practical in current high-fidelity simulations for jets, given that the likelihood is typically intractable (they are implicit models). Thus, we consider Ginkgo (Cranmer, Kyle et al., 2021; Cranmer et al., 2019b): a semi-realistic generative model for jets with a tractable joint likelihood and captures essential ingredients of parton shower generators in full physics simulations. In particular, Ginkgo was designed to enable implementations of probabilistic programming, differentiable programming, dynamic programming and variational inference.Within the analogy between jets and NLP, Ginkgo can be considered as ground-truth parse trees with a known language model.
1.4 Related work
1.4.1 Jet clustering
@@
For each jet produced at the LHC, there is an inference task on the latent hierarchy that typically involves 10 to 100 particles (leaves). Though this is a relatively small number of elements, exhaustive solutions are intractable, and current exact methods, e.g., Greenberg et al. (2020; 2021), have limited scalability. The industry standard uses agglomerative clustering techniques, which are greedy and based on heuristics (Cacciari et al., 2008; Catani et al., 1993; Dokshitzer et al., 1997; Ellis & Soper, 1993), typically finding low-quality hierarchical clusterings. Regarding likelihood-based clustering (applied to Ginkgo datasets), previous work Greenberg et al. (2020) introduced a classical data structure and dynamic programming algorithm (the cluster trellis) that exactly finds the marginal likelihood over the space of configurations and the maximum likelihood hierarchy. Also, an A* search algorithm combined with a trellis data structure that finds the exact maximum likelihood hierarchy was introduced in  Greenberg et al. (2021). Finally,  Cranmer et al. (2022) pairs Ginkgo with the cluster trellis (Greenberg et al., 2020), to use the marginal likelihood to directly characterize the discrimination power of the optimal classifier (J. Stuart & Arnold, 1994; Cranmer & Plehn, 2007) as well as to compute the exact maximum likelihood estimate for the simulator’s parameters. While these works provide exact algorithms that extend the reach of brute force methods, they have an exponential space and time complexity, becoming intractable for datasets with as few as 15 leaves. For this reason, Greenberg et al. (2020; 2021) also provide approximate solutions at the cost of finding lower-quality hierarchies.
1.4.2 Bayesian inference
@@
A recent body of research has melded variational inference (VI) and sequential search. These connections are realized through the development of a variational family for hidden Markov models, employing Sequential Monte Carlo (Smc) as the marginal likelihood estimator (Maddison et al., 2017; Naesseth et al., 2018; Le et al., 2018; Moretti et al., 2019; 2020; 2021). Within the field of Bayesian phylogenetics (the study of evolutionary histories), various methods have been proposed for inference on tree structures. Common approaches include local search algorithms like random-walk Mcmc (Ronquist et al., 2012) and sequential search algorithms like Combinatorial Sequential Monte Carlo (Csmc(Bouchard-Côté et al., 2012; Wang et al., 2015). Mcmc methods also handle model learning.  Dinh et al. (2017) proposes ppHmc which extends Hamiltonian Monte Carlo to phylogenies. Evaluating the likelihood term in Mcmc acceptance ratios can be challenging. As a workaround, particle Mcmc (Pmcmc) algorithms use Smc to estimate the marginal likelihood and define Mcmc proposals for parameter learning (Wang & Wang, 2020).
@@
Pseudo-marginal methods are a class of statistical techniques used to approximate difficult-to-compute probabilities, typically by introducing auxiliary random variables to form an unbiased estimate of the target probability (Andrieu & Roberts, 2009).  Beaumont (2003) introduced a method in genetics to sample genealogies in a fully Bayesian framework. Tran et al. (2016) utilizes pseudo-marginal methods to perform variational Bayesian inference with an intractable likelihood. Our work is a synthesis of Wang et al. (2015) and Moretti et al. (2021) in that we introduce a variational approximation on topologies using Smc and a VI framework to learn parameters.
1.5 Contributions of this Paper:
@@
    We expand upon the Csmc technique introduced by Wang et al. (2015) and the Ncsmc method from Moretti et al. (2021) to introduce a Combinatorial Sequential Monte Carlo framework for inferring hierarchical clusterings for jets. The resulting estimators are unbiased and consistent. To the best of our knowledge, this marks the first adaptation of Smc methods to jet reconstruction in particle physics.
@@
    We leverage the resulting Smc estimators to develop two approximate posteriors on jet hierarchies and correspondingly two VI methods for parameter learning. We illustrate the effectiveness of both methods through experiments using data generated with Ginkgo (Cranmer, Kyle et al., 2021), highlighting superior speed and accuracy across various tasks.
@@
    We show how partial states and re-sampled indices generated by Smc can be interpreted as auxiliary random variables within a pseudo-marginal framework, thus establishing connections between variational pseudo-marginal methods and Vsmc (Naesseth et al., 2018; Moretti et al., 2021).
@@ 2 Background
Section 2.1 provides an overview of the Ginkgo generative model for jet physics (Cranmer, Kyle et al., 2021). Section 2.3.1 summarizes the approximate inference techniques this work builds upon.
@@ 2.1.1 Generative process
In Ginkgo the generative process begins with the splitting of a parent (root) node with invariant mass squared $t_{P}$ into two children, as shown schematically in Fig. 2 (left). The process is characterized by a cutoff mass squared $t_{cut}$, and a rate parameter $\lambda$ for the exponential distribution governing the decay. During generation, as long as the invariant mass squared of a node exceeds the cutoff value ($t_{P}>t_{cut}$), that node is promoted to be a parent and the algorithm recursively splits it. The squared masses of each new left (L) and right (R) child nodes ($t_{L}$, $t_{R}$) are obtained from sampling from the exponential distribution $f(t|\lambda,t_{P})$ defined in Eq. 1, with parameters specific to each child ($L$, $R$). Finally, once $t_{L}$ and $t_{R}$ are sampled, the corresponding energy-momentum vectors ($z_{L}$, $z_{R}$) for the ($L$, $R$) nodes are derived from $t_{P}$, $t_{L}$ and $t_{R}$ following energy-momentum conservation rules, i.e. applying a 2-body particle decay (see  (Cranmer, Kyle et al., 2021; Cranmer et al., 2019b) for more details). Next, we specify the exponential distribution
@@
$t_{P}=t(z_{P})>t_{cut}$$t_{L}\sim f(\cdot|\lambda,t_{P})$$t_{R}\sim f(\cdot|\lambda,(\sqrt{t_{P}}-\sqrt{t_{L}})^{2})$LR
@@
$t_{P}=t(z_{L}+z_{R})$$\ell(t_{L},\lambda,t_{cut},t_{P}^{L},t_{P})$$=f(t|\lambda,t_{P})$$\ell(t_{R},\lambda,t_{cut},t_{P}^{R},t_{P})$$=f(t|\lambda,(\sqrt{t_{P}}-\sqrt{t_{L}})^{2})$$t_{L}=t(z_{L})$$t_{R}=t(z_{R})$
@@
Figure 2: Illustration of the Ginkgo generative and reconstruction processes. (a) Ginkgo starts with a parent node characterized by a 4-vector $z_{p}=(E,\vec{p})$ and invariant mass squared $t_{P}=E^{2}-|\vec{p}|^{2}$. If $t_{P}$ is greater than the cut off value $t_{cut}$, then the parent node splits (we have a particle decay). The left and right nodes invariant mass squared ($t_{L}$, $t_{R}$) are sampled from a truncated exponential distribution defined in Eq. 1. (b) The splitting likelihood reconstruction process of a node, defined in Eq. 4 begins with two child nodes $L$ and $R$ along with their respective 4-vectors $z_{L}$ and $z_{R}$. The 4-vector for the parent node $P$ is calculated as $z_{P}=z_{L}+z_{R}$ and then $t_{P}=t(z_{P})$. Next, we obtain $t_{L}=t(z_{L})$, $t_{R}=t(z_{R})$ and define $t_{P}^{L}=t_{P}$, and $t_{P}^{R}=(\sqrt{t_{P}}-\sqrt{t_{L}})^{2}$. Finally, the left splitting term likelihood $\ell(t_{L},\lambda,t_{cut},t_{P}^{L},t_{P})$ and the right one $\ell(t_{R},\lambda,t_{cut},t_{P}^{R},t_{P})$ are evaluated.